# Staying on the Manifold: Geometry-Aware Noise Injection

### v1

@@ 5 Experimental validation
We test our hypothesis on a range of manifolds visualised in Figure 5. We generate $N=200$ training points on each manifold and train an overparametrised 3-layer neural network with 64 nodes per layer to learn a specific function for each manifold. We train for $500$ epochs using a learning rate of $10^{-3}$ with a MSE objective. For the DeformedSphere we only use $N=40$ and a learning rate of $0.005$ for computational speed-up. For each training step, we add either ambient space noise, tangential noise, geodesic noise or Brownian motion noise and compare to a baseline network trained without adding input noise. We treat the noise covariance $\sigma^{2}$ as a hyperparameter, and, in the Brownian motion setting, interpret it as the total time of the process, i.e. $T=\sigma^{2}$. We provide the average error per method relative to the baseline’s MSE in Table 1 with uncertainties given by the standard error of the mean computed from $5$ independent runs. We provide computations for the geodesic equations and Brownian motion along with the target functions for the parametrised manifolds in Appendix A.
Our results show that geometry-aware noise injection provides advantages to ambient space noise on complex manifolds. In particular, geodesic and Brownian motion noise not only yield lower errors on ”wigglier” geometries - such as the SwissRoll and DeformedSphere - but they also exhibit greater robustness to the noise intensity hyperparameter (Figure 6). This indicates that geometric approaches can both improve generalisation and reduce sensitivity to hyperparameter choices.
@@
At the same time, performance never significantly deteriorates when using any noise strategy, compared to the baseline trained without noise (Table 1). For some manifolds simple ambient Gaussian noise can suffice: we observe this particularly for manifolds of which only a small part is problematic, such as the Bead (the fat torus). Here, Gaussian noise only leads to misleading samples near the genus. Since the surface area of the genus is proportionally small, the overall error remains low. The SwissRoll, on the other hand, is sensitive to Gaussian noise everywhere, and hence our methods work better. For completeness, we report results across all tested manifolds, even when geometry-aware strategies do not provide measurable gains.
@@ 6 Related Work
A recent work [3] surveys classical perspectives and modern advances for how noise injection influences learning . Bishop’s analysis [4] of Gaussian input noise is restricted to considering the Gauss-Newton part of the Hessian with the argument that the non-linear modelling error (NME) vanishes at an optimum [4]. The implications of this were recently discussed in the paper [21], with the conclusion that the neglected Hessian component can play a crucial role in shaping the geometry of the loss surface.
Instead of assuming that the input points live on a manifold, we can also enforce that the parameters of the model belong to a manifold. A previous work [5] analyses the impact of adding Gaussian noise to weights of a parametric model . Other works [22, 23] study orthogonal regularisers on the weight matrices, promoting the columns to be orthonormal. These constraints restrict the parameter space to the Stiefel or Grassmann manifolds, which improves numerical stability. This line of work highlights that geometry can be injected not only through noise in the input space but also by shaping the structure of the model’s parameters. Another approach to noise injection considers adding structured noise to the gradient during training with gradient-based optimisers for improved generalisation [24, 25].
Figure 6: Test loss as a function of noise intensity $\sigma^{2}$ for different noise injection strategies. The geometry-aware noise strategies that stay on the manifold, i.e. geodesic noise and Brownian motion noise, show greater robustness to the noise intensity compared to ambient or tangential noise. All methods perform at least as well as training without noise (dashed line).
In the context of Riemannian representation learning, adding noise according to the structure of the manifold stabilises results in the recent paper [26]. This approach replaces the traditional encoder–decoder setup with a Riemannian generative decoder. It directly optimises manifold-valued latent variables via a Riemannian optimiser, thereby avoiding the difficulties of approximating densities on complex manifolds. By enforcing the manifold structure during training, the learnt latent representations remain aligned with the intrinsic geometry of the data, leading to more interpretable models and stable training dynamics.
The tangent plane of a data manifold is approximated through singular value decomposition and used for sampling points in alignment with the data’s structure in a recent work [27]. This resembles our tangent space projected noise. For the methodology of the geodesic noise, a closely related idea has been explored in the context of Riemannian Laplace approximations for Bayesian inference in deep neural networks [28, 29].
@@ 7 Conclusion
We have established several geometry-aware noise injection strategies and demonstrated their need through theoretical and experimental contributions. Further, we have shown their qualities and shortcomings. In particular, we find that while ambient Gaussian noise is simple and may improve performance on nearly Euclidean manifolds, it falls short on more curved or ”wiggly” manifolds, where geodesic and Brownian motion noise provide clear advantages. These geometry-aware strategies not only improve generalisation, but are also more robust to the noise intensity. The latter reduces the burden of hyperparameter tuning. Additionally, we proposed a framework for deforming parametrised manifolds to arbitrary manifolds, which extends the use of our methods beyond standard benchmark geometries. However, we remark that this added flexibility currently comes with increased computational cost.
@@
We expect our approach to extend to higher dimensions. As explained in the introduction, a large difference between the dimensions of the ambient space and the data manifold could lead to more dramatic results as Gaussian noise samples will with high probability be normal to the manifold. However, this increases the computational complexity.
One direction we hope to explore is to consider manifolds of which we have no explicit parametrisation. One approach is to approximate the data manifold with a generative model, e.g. as previously done with variational autoencoders [30] or explore flow matching techniques as established in Section 4. Another promising direction for geometry-aware noise for point clouds is topological data analysis [31]. Lastly, an idea is to approximate the tangent plane of a data manifold as in [27] and use it for generating geodesics.
@@ Acknowledgments
This work was supported by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516), and by the DFF Sapere Aude Starting Grant “GADL”.
Appendix A Computations on the example manifolds
A.1 Biconcave disc
The biconcave disc yields an approximation of human erythrocytes, as shown in [9]. Letting $r=\sqrt{u^{2}+v^{2}},$ and let $a,b,c,d$ be parameters, then the height function for the upper half is given by
@@ A.1 Biconcave disc
Here, $d$ describes the diameter, $a$ the height at the centre, $b$ the height of the highest point, and $c$ the flatness in the centre. A parametrisation of the upper half of this surface of rotation is given by
@@
A.1.1 Tangential noise on the biconcave disc
@@
An standard computation shows that
@@
A.1.2 Geodesics on the biconcave disc
@@
A.1.3 Brownian motion on the biconcave disc
For the Brownian motion, we yield
@@ A.1.3 Brownian motion on the biconcave disc
$+\frac{1}{\sqrt{1+\frac{\partial z}{\partial r}^{2}}}dB(t)_{1}$
@@
A.2 Spheroids
We consider manifolds which are squeezed spheres. For $a,c\in\mathbb{R}^{+},$ consider the parametrisation $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ given by
@@ A.2 Spheroids
A.2.1 Tangential noise on the spheroid
@@ A.2.2 Geodesics on the spheroid
A.2.3 Brownian motion on the spheroid
@@
A.3 Tori
We also investigate different tori, some more like onion rings, others more like beads. For coefficients $a,c\in\mathbb{R}^{+},$ they can be parametrised by $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3},$
@@ A.3 Tori
A.3.1 Tangential noise on tori
@@
A.3.2 Geodesic noise on tori
@@
$\ddot{\alpha}_{1}(t)$ $=$ $\frac{(a+c\sin{\alpha_{1}(t)})\cos{\alpha_{1}(t)}}{c}\cdot\dot{\alpha}_{2}(t)^{2}$
@@
A.3.3 Brownian motion on tori
@@
$du_{k}(t)$ $=$ $\begin{bmatrix}\frac{\cos{u}}{2c(a+c\sin{u})}\\ 0\end{bmatrix}_{k}dt,$
@@
Appendix B Implementation details
B.1 Geodesic noise
To simplify our computations, instead of sampling a vector ${\boldsymbol{\epsilon}}_{\top}\sim\mathcal{N}\left(0,\sigma^{2}\mathbf{P}\right)$ in the tangent space, we can also sample a vector $\tilde{{\boldsymbol{\epsilon}}}$ in the parameter space $\mathbb{R}^{d}$ from an adjusted distribution. In the following assume that $\boldsymbol{u}\in\mathbb{R}^{d},\ \ X(\boldsymbol{u})={\boldsymbol{x}}\in\mathcal{M},$ where $X$ is a smooth parametrisation of a regular manifold $\mathcal{M}.$ As previously described, the Jacobian transforms vectors in the parameter space to the tangent space, i.e. for a vector ${\boldsymbol{\epsilon}}\in T_{\boldsymbol{u}}\mathbb{R}^{d},$ we have that
${\boldsymbol{\epsilon}}=\mathbf{J}_{X}\tilde{{\boldsymbol{\epsilon}}}\in T_{\boldsymbol{x}}\mathcal{M}$
For the inverse relation, we obtain
Consequently, if
${\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{D}),$
then for its tangential component it holds that
${\boldsymbol{\epsilon}}_{\top}\sim\mathcal{N}(0,\sigma^{2}\mathbf{P}),$
and for the pull-back it holds that
We now can find the curve $\alpha:\mathbb{R}\rightarrow\mathbb{R}^{d}$ such that
Our new sample point is then
[…]
This method is equivalent to the one described in Subsection 3.2. For simplicity, we ignore the injectivity radius of the domain of the exponential map. This is not a problem since we do not require injectivity for our purposes and the manifolds we consider are complete.
B.2 Functions on the manifolds
For the Sphere, SqueezedSphere and DeformedSphere, we select the target function as:
$y=v$
i.e. the second local coordinate.
For the Bead we select the target function as:
$y=\sin v$
i.e. a periodic function of the second local coordinate.
For the OnionRing we select the target function as:
$y=100\cdot c\cdot\cos u=100\cdot z$
i.e. the scaled height of the manifold.
For the SwissRoll we select the target function as:
$y=u$
i.e. the first local coordinate, which is a linearly increasing function along the roll.
Appendix C Details on manifold deformations

### v2

@@ 5.1 Parameterised Manifolds in $\mathbb{R}^{3}$
We first test our hypothesis on a range of parameterised manifolds in $\mathbb{R}^{3}$. We generate $N=200$ training points on each manifold and train an overparameterised 3-layer neural network with 64 nodes per layer to learn a specific function for each manifold. We train for $500$ epochs using a learning rate of $10^{-3}$ with a MSE objective. For the DeformedSphere we only use $N=40$ and a learning rate of $0.005$ for computational speed-up. For each training step, we add either ambient space noise, tangential noise, geodesic noise or Brownian motion noise and compare to a baseline network trained without adding input noise. We treat the noise covariance $\sigma^{2}$ as a hyperparameter, and, in the Brownian motion setting, interpret it as the total time of the process, i.e. $T=\sigma^{2}$. We provide the average error per strategy relative to the baseline’s MSE in Table 1 with uncertainties given by the standard error of the mean computed from $5$ independent runs. We provide computations for the geodesic equation and Brownian motion along with the target function for each manifold in Appendix B.
@@
Our results show that geometry-aware noise injection provides advantages to ambient space noise on complex manifolds. In particular, geodesic and Brownian motion noise yield lower errors on ”wigglier” geometries, such as the SwissRoll, and they also exhibit greater robustness to the noise intensity hyperparameter (Figure 5). This indicates that geometric approaches can both improve generalisation and reduce sensitivity to hyperparameter choices. At the same time, performance rarely significantly deteriorates when using any noise strategy, compared to the baseline trained without noise (Table 1). For some manifolds, simple ambient Gaussian noise can suffice, particularly for those of which only a small part is problematic, such as the Bead (the fat torus). Here, Gaussian noise only leads to misleading samples near the genus. Since the surface area of the genus is proportionally small, the overall error remains low. The SwissRoll, on the other hand, is sensitive to Gaussian noise everywhere, and our strategies work better. For completeness, we report results across all manifolds, even when geometry-aware strategies do not provide measurable gains.
Figure 5: Test loss on the SwissRoll as a function of noise intensity $\sigma^{2}$ for different noise injection strategies. The geometry-aware noise strategies that stay on the manifold, i.e. geodesic noise and Brownian motion noise, show greater robustness to the noise intensity compared to ambient or tangential noise. Our strategies perform at least as well as training without noise (dashed line).
Which is the better strategy?
Though both geodesic noise and Brownian motion noise perform equally well under certain conditions, Brownian motion noise is computed more efficiently than geodesic noise, which requires solving the exponential map with high precision. Due to the stochastic nature of Brownian motion, it is less affected by the resolution of the time discretisation which allows for speeding up the sampling process. For these reasons, we restrict further analyses to only consider our geometry-aware Brownian motion strategy.
5.2 MNIST
Figure 6: Four different augmentations of a specific image of a ’0’ under ambient Gaussian noise (top) and geometry-aware Brownian motion noise (bottom) for a relatively large noise intensity of $\sigma=1$. The geometry-aware samples resemble digits and thus stay on the manifold which is not the case for the ambient noise samples.
Figure 7: An example image of a ’0’ (left) along with the pixel-wise standard deviation over $100$ augmentations of it, using ambient Gaussian noise (middle) and intrinsic Brownian motion (right). Where ambient noise is somewhat uninformative, the geometry-aware noise targets natural variations on the edges of the digit.
We now turn our attention to a higher-dimensional example using image data, where the manifold must be approximated. One common approach to approximate such a manifold is using autoencoders [arvanitidis2017latent]. An autoencoder uses an encoder-decoder structure to reconstruct input data samples with minimum reconstruction error. As such, an autoencoder has an inherent latent space $\mathcal{Z}\subseteq\mathbb{R}^{d}$ in which we can represent the data samples using the encoder, i.e. $\boldsymbol{z}=f_{e}\left({\boldsymbol{x}}\right)$. The reconstruction is obtained by decoding the latent representation to a point on the approximate data manifold $\widetilde{\mathcal{X}}\subseteq\mathbb{R}^{D}$, i.e. $\tilde{{\boldsymbol{x}}}=f_{d}\left(\boldsymbol{z}\right)$. We can therefore think of the latent space $\mathcal{Z}$ as the parameter space of the approximate data fold $\widetilde{\mathcal{X}}$ with the decoder serving as the chart (similar to Figure 3). Typically, $d\ll D$ which makes it favourable for doing manipulations of the data and defining the pullback metric of the approximated data manifold in the latent space allows us to apply our geometry-aware noise injection strategies on approximated data manifolds.
We test our hypothesis using intrinsic Brownian motion on the approximated image manifold, specifically on the MNIST digits dataset. First, we train an autoencoder on the full training set. Next, we train a 1-layer MLP classifier with $1024$ hidden units with various levels of Brownian motion noise added to the data during training. We compare to adding Gaussian noise in the ambient image space. The MNIST dataset consists of $60,000$ samples and covers the image manifold of digits well, for which reason we test our strategy in settings of subsampling the training data to $1\%,10\%$ or $50\%$ of the dataset. We do so to examine highly overparameterised settings where data augmentation is expected to improve the model fit. See Appendix C for experimental details.
In Figure 6, we show different augmentations with ambient noise and intrinsic Brownian motion noise for an example of a ’0’. While our geometry-aware approach generates digit-looking images, the underlying signal is hard to recognise in the case of Gaussian noise. In Figure 7 we show the associated pixel-wise variations across $100$ augmented samples for each method. While the ambient noise variation is somewhat uniform, the geometry-aware samples give natural variations along the edges of the digit.
Table 2: Test accuracy on MNIST when trained on the original images (O), reconstructed images (R), original images with ambient noise (A) and reconstructed images with geometry-aware noise (BM). The header refers to the subsampling rate of the training set and uncertainties are the standard errors over 10 independent runs. We highlight the best performance per subsampling rate in bold. Green cells indicate whether training with noise is significantly better than training on the original or reconstructed images for the respective noising strategies.
$1\%$ $10\%$ $50\%$
O $0.883\pm 0.008$ $0.956\pm 0.002$ $\mathbf{0.981}\pm\mathbf{0.001}$
A $0.883\pm 0.008$ $\mathbf{0.965}\pm\mathbf{0.001}$ $\mathbf{0.981}\pm\mathbf{0.001}$
R $0.877\pm 0.005$ $0.943\pm 0.002$ $0.967\pm 0.002$
BM $\mathbf{0.896}\pm\mathbf{0.008}$ $0.959\pm 0.002$ $0.971\pm 0.001$
In the most overparameterised setting using $1\%$ of the data, our geometry-aware noise injection strategy shows improved performance over learning without noise and learning with ambient Gaussian noise (see Table 2). In this setting, we additionally see that increasing the noise intensity of the ambient noise deteriorates the classifier’s performance, while the trend is opposite for the geometry-aware noise strategy (Figure 9). It is worth noting the performance gap of approximately $0.6\%$ when trained on the original images compared to the reconstructed images, yet we highlight that the geometry-aware noise eventually surpass this gap.
When the classifier is trained on larger amounts of the training set, the performance gap between training on the original and reconstructed images grows, resulting in the geometry-aware strategy not improving over training without noise on the original images. Yet, we note that our strategy performs consistently better than the classifier trained directly on the reconstructed images. We therefore expect the strategy to work well if lowering the autoencoder’s approximation error, i.e. learning a better approximation of the data manifold. We remark that learning a perfect approximation of the data manifold is not the aim of this paper.
One potential limitation of the geometry-aware strategy is that the augmented samples might resemble other digits than the label associated with the original sample. This is due to the fact that the strategy does not have information about the digit labels from the decoder itself. If the intrinsic Brownian motion crosses the label boundary, it can negatively impact the classifier performance due to label noise. For intuition, see the example of transitioning from a ’$4$’ to a ’$9$’ in Figure 8. We considered solving this potential issue by also pulling back information from the classifier activations to the latent space, however initial experiments revealed no significant performance gain. This could be due to the fact that the augmented samples already lie along the label border, giving a stronger and more robust classifier.
Figure 8: Time slices of the geometry-aware Brownian motion starting from an image of a ’$4$’ at time $t=0$. As the decoder is unaware of the digit labels, the Brownian motion can cross the label boundary, resulting in the augmentation at time $t=1$ resembling a ’$9$’.
@@ 6 Related Work
A recent work [ferianc2024navigatingnoisestudynoise] surveys classical perspectives and modern advances for how noise injection influences learning. Instead of assuming that the input points live on a manifold, we can also enforce that the parameters of the model belong to a manifold. A previous work [an1996effects] analyses the impact of adding Gaussian noise to weights of a parametric model. Other works [massart2022orthogonal, massion2024minimizers] study orthogonal regularisers on the weight matrices, promoting the columns to be orthonormal. These constraints restrict the parameter space to the Stiefel or Grassmann manifolds, which improves numerical stability. This line of work highlights that geometry can be injected not only through noise in the input space but also by shaping the structure of the model’s parameters. Other works inject noise to the gradient during training with gradient-based optimisers for improved generalisation [orvieto2022anticorrelated, orvieto2023explicit].
In the context of Riemannian representation learning, adding noise according to the structure of the manifold stabilises results in the recent paper [bjerregaard2025riemannian]. This approach replaces the traditional encoder–decoder setup with a Riemannian generative decoder. It directly optimises manifold-valued latent variables via a Riemannian optimiser, thereby avoiding the difficulties of approximating densities on complex manifolds. By enforcing the manifold structure during training, the learnt latent representations remain aligned with the intrinsic geometry of the data, leading to more interpretable models and stable training dynamics.
In a recent work [kaufman2024first], the tangent plane of a data manifold is approximated through singular value decomposition and used for sampling points in alignment with the data structure, similar to our tangent space noise. For the methodology of the geodesic noise, a related idea has been explored in the context of Riemannian Laplace approximations for Bayesian inference in deep neural networks [bergamin2023riemannianlaplaceapproximationsbayesian, da2025geometric].
Figure 9: Test accuracy of an overparameterised 1-layer MLP trained on the $1\%$ subsampled MNIST dataset using ambient noise and our proposed geometry-aware Brownian motion strategy. For ambient noise, the model performance deteriorates, while geometry-aware Brownian motion improves generalisation.
@@ 7 Conclusion
We have established several geometry-aware noise injection strategies and demonstrated their need through theoretical and experimental contributions. Further, we have shown their qualities and shortcomings. In particular, we find that while ambient Gaussian noise is simple and may improve performance on nearly Euclidean manifolds, it falls short on more curved or ”wiggly” manifolds, where geodesic and Brownian motion noise provide clear advantages. These geometry-aware strategies not only improve generalisation, but are also more robust to the noise intensity with the latter reducing the burden of hyperparameter tuning. We proposed a framework for deforming parameterised manifolds to arbitrary manifolds, which extends the use of our strategies beyond standard benchmark geometries. However, we remark that this added flexibility currently comes with increased computational cost. Lastly, we showed an how to apply our techniques to higher-dimensional manifolds approximated by an autoencoder.
@@
Though our results in the high-dimensional setting of image data did not give strictly better performance, we attributed the performance gap to the quality of the manifold approximation, thus future work involves learning a better approximator of the manifold using e.g. flow matching as established in Section 4. We expect a large difference between the dimensions of the ambient space and the data manifold to lead to more dramatic results, as Gaussian noise samples will with high probability be normal to the manifold. Thus, studying the relation between the ambient space dimensionality, data manifold dimensionality and the classifier performance is of interest.
@@ Acknowledgments
This work was supported by Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516), and by the DFF Sapere Aude Starting Grant “GADL”. The Otto Mønsted Fond provided generous support for the authors’ travel. We also want to thank the reviewers for the helpful feedback.
Appendix A Implementation Details
A.1 Geodesic Noise
To simplify our computations, instead of sampling a vector ${\boldsymbol{\epsilon}}_{\top}\sim\mathcal{N}\left(0,\sigma^{2}\mathbf{P}\right)$ in the tangent space, we can also sample a vector $\tilde{{\boldsymbol{\epsilon}}}$ in the parameter space $\mathbb{R}^{d}$ from an adjusted distribution. In the following assume that $\boldsymbol{u}\in\mathbb{R}^{d},\ \ X(\boldsymbol{u})={\boldsymbol{x}}\in\mathcal{M},$ where $X$ is a smooth parameterisation of a regular manifold $\mathcal{M}.$ As previously described, the Jacobian transforms vectors in the parameter space to the tangent space, i.e. for a vector ${\boldsymbol{\epsilon}}\in T_{\boldsymbol{u}}\mathbb{R}^{d},$ we have that
${\boldsymbol{\epsilon}}=\mathbf{J}_{X}\tilde{{\boldsymbol{\epsilon}}}\in T_{\boldsymbol{x}}\mathcal{M}.$
For the inverse relation, we obtain
Consequently, if
${\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{D}),$
then for its tangential component it holds that
${\boldsymbol{\epsilon}}_{\top}\sim\mathcal{N}(0,\sigma^{2}\mathbf{P}),$
and for the pullback it holds that
This allows us to find the curve $\alpha:\mathbb{R}\rightarrow\mathbb{R}^{d}$ such that
Our new sample point is then
[…]
This strategy is equivalent to the one described in Subsection 3.2. For simplicity, we ignore the injectivity radius of the domain of the exponential map – this is not a problem since we do not require injectivity for our purposes, and the manifolds we consider are complete.
A.2 Functions on the Manifolds
For the Sphere, SqueezedSphere and DeformedSphere, we select the target function as
$y=v,$
that is, the second local coordinate.
For the Bead we select the target function as
$y=\sin v$
which is a periodic function of the second local coordinate.
For the OnionRing we select the target function as:
$y=100\cdot c\cdot\cos u=100\cdot z,$
which is the scaled height of the manifold.
For the SwissRoll we select the target function as:
$y=u,$
namely the first local coordinate.
Appendix B Manifold Computations
B.1 Biconcave disc
The biconcave disc yields an approximation of human erythrocytes, as shown in [kuchel2021surface]. Letting $r=\sqrt{u^{2}+v^{2}},$ and let $a,b,c,d$ be parameters, then the height function for the upper half is given by
@@ B.1 Biconcave disc
Here, $d$ describes the diameter, $a$ the height at the centre, $b$ the height of the highest point, and $c$ the flatness in the centre. A parameterisation of the upper half of this surface of rotation is given by
@@
B.1.1 Tangential noise on the biconcave disc
@@
A standard computation shows that
@@
B.1.2 Geodesics on the biconcave disc
@@
B.1.3 Brownian motion on the biconcave disc
For the Brownian motion, we have that
@@ B.1.3 Brownian motion on the biconcave disc
$+\frac{1}{\sqrt{1+\frac{\partial z}{\partial r}^{2}}}dB(t)_{1},$
@@
B.2 Spheroids
We consider manifolds which are squeezed spheres. For $a,c\in\mathbb{R}^{+},$ consider the parameterisation $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ given by
@@ B.2 Spheroids
B.2.1 Tangential noise on the spheroid
@@ B.2.2 Geodesics on the spheroid
B.2.3 Brownian motion on the spheroid
@@
B.3 Tori
We also investigate different tori, some more like onion rings, others more like beads. For coefficients $a,c\in\mathbb{R}^{+},$ they can be parameterised by $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3},$
@@ B.3 Tori
B.3.1 Tangential noise on tori
@@
B.3.2 Geodesic noise on tori
@@
$\ddot{\alpha}_{1}(t)$ $=$ $\frac{(a+c\sin{\alpha_{1}(t)})\cos{\alpha_{1}(t)}}{c}\cdot\dot{\alpha}_{2}(t)^{2},$
@@
Figure 10: Test accuracy of an overparameterised 1-layer MLP trained on different subsampled versions of the MNIST dataset using ambient noise and our proposed geometry-aware Brownian motion strategy.
B.3.3 Brownian motion on tori
@@
$du_{k}(t)$ $=$ $\begin{bmatrix}\frac{\cos{u}}{2c(a+c\sin{u})}\\ 0\end{bmatrix}_{k}dt$
@@
Appendix C MNIST Experiment
We first trained an autoencoder on the full training dataset. Both the encoder and decoder were defined as convolutional neural networks with softplus activation functions and a $d=16$ dimensional latent space. For stability, we choose the output function of the decoder to be the hyperbolic tangent. Since the hyperbolic tangent maps the real line to (1,1)(-1,1), and we considered MNIST images normalized to the pixel range of $[0,1]$, we convert the decoder output to lie in the same pixel interval. We train the autoencoder with the MSE loss objective using a batch size of $64$, a learning rate of $0.01$ and weight decay of $3\cdot 10^{-5}$ for 100 epochs using Adam and a cosine learning rate schedule acting every epoch.
Next, we train a classifier to distinguish the MNIST digits using different versions of the images: 1) the original images, 2) the reconstructed images using the autoencoder, 3) the images with ambient Gaussian noise and 4) the images based on geometry-aware Brownian motion in the latent space. For Gaussian noise in the ambient space, we clip the pixel-values to the $[0,1]$ range. We do so using either $1\%$, $10\%$, $50\%$ and $100\%$ of the training dataset. We define the classifier as a $1$-layer MLP with $1024$ hidden units using ReLU as the activation function. We use a learning rate of $0.001$ and train the classifier until convergence for 100 epochs with Adam and a cosine learning rate scheduler acting every epoch. We use the negative log-likelihood as the training objective. We repeat the experiment for noise intensities $\sigma$ log-spaced between $10^{-4}$ and $1$.
All training was repeated for 10 different random initialisations of both the autoencoder and the classifier. We compute the mean accuracy on the test set for each noise intensity along with the standard error of the mean over these 10 runs. We show the results for all settings in Figure 10. In Table 2 and 3 we show the best test accuracy (i.e. for the best noise intensity) when training the classifier on all subsampled versions.
Table 3: Test set accuracy for the best performing classifiers trained on the full training dataset.
$100\%$
Or $0.986\pm 0.001$
A $\mathbf{0.986\pm 0.001}$
Re $0.972\pm 0.002$
BM $0.976\pm 0.001$
Appendix D Manifold Deformation